Rose's garden is in the shape of a trapezoid. If the height of the trapezoid is 16m, one base is 20m, and the area is 224m second power, find the length of the other base.
Answer by checkley71(8403)

(b1+b2)/2*h=area
(20+b2)/2*16=224
(20+b2)*8=224
160+8*b2=224
8b2=224-160
8b2=64
b2=64/8
b2=8m is the second base.
proof:
(20+8)/2*16=224
28/2*16=224
14*16=224
224=224